In planning her retirement, Callie Daniels deposits some money at 2.5% interest, with twice as much deposited at 3%. Find the amount deposited a each rate if the total annual interest income is $850.
Answer by mananth(16946)

In planning her retirement, Callie Daniels deposits some money at 2.5% interest, with twice as much deposited at 3%. Find the amount deposited a each rate if the total annual interest income is $850.
..
@ 2.5 % --------------- $x
@3%--------------------$2x
...
Interest each year = $850
..
2.5x+3*2x=850*100
8.5x=85,000
/8.5
x=85,000/8.5
x=$10,000 ----- investment @ 2.5%
2x=10,000*2=$20,000 -------- investment @ 3%
...
m.ananth@hotmail.ca